[Brahms-l] Shift report 20050301 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Mar 01 2005 - 16:03:27 EST
Shifter leader was JHL
Shifters were SB


1. Beam for 6 hours with ZDC~16K-3K/sec
   8:15am-1pm, 2:15pm-

2. Data (from run13543 - 13548)

   FS 8deg. 2442A
   MRS 40deg. 700B

   Scale down factors have been optimized for
   trigger2 with DAQ live ~60-70%. 
   (160K Trigger2/hour at the beginning of store)

3. No new detector/DAQ issues
~
~


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 1 16:03:44 2005

This archive was generated by hypermail 2.1.8 : Tue Mar 01 2005 - 16:04:04 EST